# Fuzzy Information Evolution with Three-Way Decision in Social Network Group Decision-Making

### v1

@@ V-G1 Brain-Computer Interface (BCI) systems
BCI systems often involve non-invasive brain signals such as electroencephalography (EEG) and functional near-infrared spectroscopy (fNIRS). These signals usually have high variability, strong individual differences, and incomplete information. Traditional classification methods require making clear decisions for all states. However, in real-world applications, the user’s cognitive state may be “uncertain” or “hesitant.” Introducing the 3WD theory allows the brain signal recognition system to produce three types of outputs: “accept,” “reject,” and “defer.” This keeps an intermediate state for handling vague information and helps avoid incorrect triggering. In addition, the mechanism of dynamically adjusting neural information channels can also inspire the design of adaptive models for individual brain region communication.
@@
V-G2 Opinion Fusion in Smart City Governance
In smart city governance, such as public policy making, intelligent traffic management, and urban risk warning, a large amount of public opinion needs to be collected and integrated. When there are differences in views, uncertain perceptions, and vague opinions, traditional opinion fusion models can easily lead to extreme results or distorted information. The proposed model allows individual opinions to be expressed in natural language and handles uncertain areas through deferred decisions. This makes the opinion fusion process more understandable and more tolerant of ambiguity.
V-G3 Multi-Robot Coordination in Disaster Response and Agricultural Monitoring
In scenarios such as post-disaster search and rescue and agricultural monitoring, multi-robot systems need to make collaborative decisions based on environmental sensing, their own abilities, and task requirements. Traditional assignment strategies often require each robot to immediately accept or reject a task. However, in uncertain environments or when task complexity is unclear, this approach may lead to poor task allocation or wasted resources. By introducing the 3WD theory, robots are allowed to delay task decisions, so they can make better choices when more information becomes available. At the same time, by adjusting the communication network dynamically, robots can improve their cooperation based on task similarity or path overlap, which increases the overall efficiency and robustness of the system.
@@ VI Conclusion
This paper proposes a group opinion update model that combines the 3WD mechanism with a dynamic network update process. The introduced 3WD mechanism sets a “hesitation interval” to expand the traditional binary rule of accepting or rejecting neighbors into a three-way probabilistic rule. This allows an agent to accept information from another agent with a certain probability when their opinion difference lies in an uncertain range. In addition, to overcome the limitations of a static network structure, the model includes a dynamic update mechanism based on cognitive similarity. By setting probabilistic rules for adding or removing links, the social relationships between agents can adjust in real time as their opinions evolve. The proposed algorithm is applied to a multi-UAV cooperative decision-making scenario. By simulating the information exchange and dynamic negotiation among UAVs, we demonstrate that the mechanism is effective in practical distributed intelligent systems. The experimental comparisons further verify the advantages of the proposed algorithm.
As AI agents powered by large language models (LLMs) play an increasingly important role in group decision-making, how to combine the diverse solutions generated by different models has become a key issue. Future research can focus on conflict resolution between models, trust evaluation mechanisms, and integration strategies to improve the overall level of collaborative intelligence in the system. Also, we will add reinforcement learning or adaptive weighting based on past interactions to enable agents to evolve their decision strategies over time.

### v2

@@ V-D1 Brain-Computer Interface (BCI) systems
BCI systems often involve non-invasive brain signals such as electroencephalography (EEG) and functional near-infrared spectroscopy (fNIRS). These signals typically exhibit high variability, significant individual differences, and incomplete information. Traditional classification methods require making clear decisions for all states. However, in real-world applications, the user’s cognitive state may be “uncertain” or “hesitant.” Introducing the 3WD theory allows the brain signal recognition system to produce three types of outputs: “accept,” “reject,” and “defer.” This keeps an intermediate state for handling vague information and helps avoid incorrect triggering. In addition, the mechanism of dynamically adjusting neural information channels can also inspire the design of adaptive models for individual brain region communication.
@@
V-D2 Opinion Fusion in Smart City Governance
In smart city governance, such as public policy making, intelligent traffic management, and urban risk warning, a large amount of public opinion needs to be collected and integrated. When there are differences in views, uncertain perceptions, and vague opinions, traditional opinion fusion models can easily lead to extreme results or distorted information. The proposed model allows individual opinions to be expressed in natural language and handles uncertain areas through deferred decisions. This makes the opinion fusion process more comprehensible and less prone to ambiguity.
V-D3 Multi-Robot Coordination in Disaster Response and Agricultural Monitoring
In scenarios such as post-disaster search and rescue and agricultural monitoring, multi-robot systems must make collaborative decisions based on environmental sensing, their own capabilities, and task requirements. Traditional assignment strategies often require each robot to accept or reject a task immediately. However, in uncertain environments or when task complexity is unclear, this approach may lead to poor task allocation or wasted resources. By introducing the 3WD theory, robots are allowed to delay task decisions, so they can make better choices when more information becomes available. At the same time, by dynamically adjusting the communication network, robots can enhance their cooperation based on task similarity or path overlap, thereby increasing the overall efficiency and robustness of the system.
@@ VI Conclusion
This paper proposes a group opinion update model that combines the 3WD mechanism with a dynamic network update process. The introduced 3WD mechanism sets a “hesitation interval” to expand the traditional binary rule of accepting or rejecting neighbors into a three-way probabilistic rule. This allows an agent to accept information from another agent with a certain probability when their opinion difference lies in an uncertain range. In addition, to overcome the limitations of a static network structure, the model includes a dynamic update mechanism based on cognitive similarity. By setting probabilistic rules for adding or removing links, the social relationships between agents can adjust in real time as their opinions evolve. The proposed algorithm is applied to a multi-UAV cooperative decision-making scenario. By simulating the information exchange and dynamic negotiation among UAVs, we demonstrate that the mechanism is effective in practical distributed intelligent systems. The experimental comparisons further verify the advantages of the algorithm.
As AI agents powered by large language models (LLMs) play an increasingly important role in group decision-making, how to combine the diverse solutions generated by different models has become a key issue. Future research can focus on conflict resolution between models, trust evaluation mechanisms, and integration strategies to improve the overall level of collaborative intelligence in the system. Additionally, we will incorporate reinforcement learning or adaptive weighting based on past interactions to enable agents to refine their decision strategies over time.